The speed of sound is 340 m/s where a tuning fork produces the second resonance position above a closed air column that is 49.8 cm in length. The frequency of the tuning fork is ___ Hz. (Give your answer with the correct number of sig digs and do not include units in your answer)


1
Expert's answer
2021-07-25T09:18:15-0400

The second resonance position corresponds to the third harmonic of the closed air column. Therefore, we can find the wavelength of the wave:


"\lambda=\frac{4}{3}L=\frac{4}{3}(0.498)=0.664\ m"

Finally, we can find the frequency of the tuning fork:


"f=\frac{v}{\lambda}=\frac{340}{0.664}=510\ Hz"